# Nearest-Neighbours Neural Network architecture for efficient sampling of statistical physics models

### v1

@@ Discussion
In this work we have introduced a new, physics-inspired 4N architecture for the approximation of the Gibbs-Boltzmann probability distribution of a spin glass system of $N$ spins. The 4N architecture is inspired by the previously introduced TwoBo [25], but it implement explicitly the locality of the physical model via a Graph Neural Network structure.
@@ Methods
The Nearest Neighbors Neural Network (4N) architecture uses a GNN structure to respect the locality of the problem, as illustrated in Fig. 1. The input, given by the local fields computed from the already fixed spins, is passed through $\ell$ layers of weights $W_{ij}^{k}$. Each weight corresponds to a directed nearest neighbor link $i\leftarrow j$ (including $i=j$) and $k=1,\dots,\ell$, therefore the network has a total number of parameters equal to $(c+1)\times\ell\times N$, i.e. linear in $N$ as long as the number of layers and the model connectivity remain finite.
@@
Autoregressive models such as 4N are based on an ordering of spins to implement Eq. (6). One can choose a simple raster scan (i.e., from left to right on one row and updating each row from top to bottom sequentially) or any other scheme of choice. In this work, we have adopted a spiral scheme, in which the spins at the edges of the system are updated first, and then one moves in spirals towards the central spins. As shown in the SI, it appears as4N performs better, for equal training conditions, when the spins are updated following a spiral scheme with respect to, for instance, a raster scheme.
@@
We use as training set of $M=2^{16}$ configurations obtained at several temperatures using a Parallel Tempering algorithm for different instances of the couplings $J_{ij}$, and for a 16x16 two-dimensional square lattice with periodic boundary conditions. Data where gathered in runs of $2^{27}$ steps by considering $2^{26}$ steps for thermalization and then taking one configuration each $2^{10}$ steps. Temperature swap moves, in which one swaps two configurations at nearby temperature, were attempted once every 32 normal Metropolis sweeps. The choice of $M=2^{16}$ was motivated based on previous studies of the same model [15].
@@
The training has been carried out maximizing the likelihood of the data, a procedure which is equivalent to minimizing the KL divergence $D_{\text{KL}}(P_{\text{GB}}\parallel P_{\text{NN}})$, with $P_{\text{GB}}$ estimated by the training set. We used an ADAM optimizer with batch size 100 for 80 epochs, early stopping with patience 10, learning rate 0.01 with exponential decay in time.
Acknowledgments
@@
We especially thank Indaco Biazzo and Giuseppe Carleo for several important discussions on their TwoBo architecture that motivated and inspired the present study. We also thank Giulio Biroli, Patrick Charbonneau, Simone Ciarella, Marylou Gabrié, Leonardo Galliano, Jeanne Trinquier, Martin Weigt and Lenka Zdeborová for useful discussions. The research has received financial support from ICSC - Italian Research Center on High-Performance Computing, Big Data, and Quantum Computing, funded by the European Union - NextGenerationEU.
S1 Probability distribution of the overlap for $L=32$
Figure S1: Additional examples of the probability distribution of the overlap, $P(q)$, for different instances of the disorder and different temperatures for the $L=32$, $N=1024$ case. Solid color lines correspond to configurations generated using 4N, while dashed black lines correspond to the $P(q)$ obtained via PT. With the increase of the system size, more layers are required in order to reproduce correctly the probability distribution of the overlap at lower temperatures, as expected by the fact that finite-size effects are less present.
@@
Figure S2: Comparison of the probability distribution of the overlap obtained using different architectures: 4N with $\ell=10$ layers (12580 parameters), shallow MADE (32640 parameters), NADE with $N_{h}=64$ (32768 parameters) and TwoBo (7709 parameters). The black dashed line is obtained using data gathered via Parallel Tempering (PT), while solid color lines are obtained using the different architectures. 4N performs comparably or better than MADE and NADE despite having less parameters. Moreover, it also performs comparably to TwoBo. Despite the latter having fewer parameters in this $N=256$ case, 4N has a better scaling when increasing the system size, $\mathcal{O}(N)$ compared to $\mathcal{O}(N^{\frac{3}{2}})$, as pointed out in the main text.
@@
Figure S3: Comparison of the probability distribution of the overlap obtained using different update sequences of the spins. The black dashed line is obtained using data gathered via Parallel Tempering (PT), while solid color lines are obtained using 4N with different sequence updates. Spiral: spins are update in a following a spiral going from the outside to the inside of the lattice; Raster: spins are updated row by row from left to right. It appears as the spiral update performs much better than the raster update for the 4N architecture in equal training conditions, something that is not observed when training, for instance, the MADE architecture. This effect could be related to the locality characteristics of the 4N architecture.

### v2

@@ Discussion
In this work we have introduced a new, physics-inspired 4N architecture for the approximation of the Gibbs-Boltzmann probability distribution of a spin glass system of $N$ spins. The 4N architecture is inspired by the previously introduced TwoBo [21], but it implements explicitly the locality of the physical model via a Graph Neural Network structure.
@@ Methods
The Nearest Neighbors Neural Network (4N) architecture uses a GNN structure to respect the locality of the problem, as illustrated in Fig. 1. The input, given by the local fields computed from the already fixed spins, is passed through $\ell$ layers of weights $W_{ij}^{k}$. Each weight corresponds to a directed nearest neighbor link $i\leftarrow j$ (including $i=j$) and $k=1,\dots,\ell$, therefore the network has a total number of parameters equal to $(c+1)\times\ell\times N$, i.e. linear in $N$ as long as the number of layers and the model connectivity remain finite. A showcase of the real wall-clock times required for the training of the network and for the generation of new configurations is shown in the SI.
@@
Autoregressive models such as 4N are based on an ordering of spins to implement Eq. (6). One can choose a simple raster scan (i.e., from left to right on one row and updating each row from top to bottom sequentially) or any other scheme of choice, see e.g. Refs. [14, 36] for a discussion. In this work, we have adopted a spiral scheme, in which the spins at the edges of the system are updated first, and then one moves in spirals towards the central spins. As shown in the SI, it appears as 4N performs better, for equal training conditions, when the spins are updated following a spiral scheme with respect to, for instance, a raster scheme.
@@
We use a training set of $M=2^{16}$ configurations obtained at several temperatures using a Parallel Tempering algorithm for different instances of the couplings $J_{ij}$, and for a 16x16 two-dimensional square lattice with open boundary conditions. Data where gathered in runs of $2^{27}$ steps by considering $2^{26}$ steps for thermalization and then taking one configuration each $2^{10}$ steps. Temperature swap moves, in which one swaps two configurations at nearby temperature, were attempted once every 32 normal Metropolis sweeps. The choice of $M=2^{16}$ was motivated based on previous studies of the same model [12].
@@
The training has been carried out maximizing the likelihood of the data, a procedure which is equivalent to minimizing the KL divergence $D_{\text{KL}}(P_{\text{GB}}\parallel P_{\text{NN}})$, with $P_{\text{GB}}$ estimated by the training set. We used an ADAM optimizer with batch size 128 for 80 epochs and a learning rate 0.001 with exponential decay in time. Training is stopped if the loss has not improved in the last 10 steps, and the lowest-loss model is considered.
Acknowledgements.
@@
We especially thank Indaco Biazzo and Giuseppe Carleo for several important discussions on their TwoBo architecture that motivated and inspired the present study. We also thank Giulio Biroli, Patrick Charbonneau, Simone Ciarella, Marylou Gabrié, Leonardo Galliano, Jeanne Trinquier, Martin Weigt and Lenka Zdeborová for useful discussions. We acknowledge support from the computational infrastructure DARIAH.IT, PON Project code PIR01_00022, National Research Council of Italy. The research has received financial support from the “National Centre for HPC, Big Data and Quantum Computing - HPC”, Project CN_00000013, CUP B83C22002940006, NRP Mission 4 Component 2 Investment 1.5, Funded by the European Union - NextGenerationEU. Author LMDB acknowledges funding from the Bando Ricerca Scientifica 2024 - Avvio alla Ricerca (D.R. No. 1179/2024) of Sapienza Università di Roma, project B83C24005280001 – MaLeDiSSi.
S1 Computational details
We now introduce a easy way to generate efficiently data using the 4N architecture. As long as the activation functions $\phi^{k}$ of the different layers of 4N are linear, all the intermediate message passing layers can be substituted with a single matrix multiplication. Indeed, let us consider, as in the main text, identity activation functions. In this case, we can rewrite the matrix implementing the $t-$th layer, $\underline{\underline{A}}^{t}$, as $\underline{\underline{A}}^{t}=\beta(\underline{\underline{J}}+\underline{\underline{I}})\odot\underline{\underline{W}}^{t}$, where $\beta$ is the inverse temperature, $\underline{\underline{J}}$ is the coupling matrix, $\underline{\underline{I}}$ is the identity matrix, and $\underline{\underline{W}}^{t}$ it the $t-$th layer weight matrix. Here, $\odot$ is the elementwise multiplication. Now we observe that (in 4N we put skip connections after all but the last layer)
$\mathbf{h}^{t}=\underline{\underline{A}}^{t}\mathbf{h}^{t-1}=\underline{\underline{A}}^{t}(\underline{\underline{A}}^{t-1}\mathbf{h}^{t-2}+h^{0})=\dots=(\underline{\underline{A}}^{t}\underline{\underline{A}}^{t-1}\dots\underline{\underline{A}}^{1}+\underline{\underline{A}}^{t}\underline{\underline{A}}^{t-1}\dots\underline{\underline{A}}^{2}+\dots+\underline{\underline{A}}^{t}\underline{\underline{A}}^{t-1}+\underline{\underline{A}}^{t})\mathbf{h}^{0}.$ (16)
Now, remembering that, if the input configuration $\mathbf{s}$ is properly masked, we have $\mathbf{h}^{0}=\underline{\underline{J}}\mathbf{s}$, it follows that
$\mathbf{h}^{t}=(\underline{\underline{A}}^{t}\underline{\underline{A}}^{t-1}\dots\underline{\underline{A}}^{1}+\underline{\underline{A}}^{t}\underline{\underline{A}}^{t-1}\dots\underline{\underline{A}}^{2}+\dots+\underline{\underline{A}}^{t}\underline{\underline{A}}^{t-1}+\underline{\underline{A}}^{t})\underline{\underline{J}}\mathbf{s},$ (17)
which can be rewritten in matrix form as $\mathbf{h}^{t}=\underline{\underline{M}}\mathbf{s}$ by defining
$\underline{\underline{M}}=(\underline{\underline{A}}^{t}\underline{\underline{A}}^{t-1}\dots\underline{\underline{A}}^{1}+\underline{\underline{A}}^{t}\underline{\underline{A}}^{t-1}\dots\underline{\underline{A}}^{2}+\dots+\underline{\underline{A}}^{t}\underline{\underline{A}}^{t-1}+\underline{\underline{A}}^{t})\underline{\underline{J}},$ (18)
and the final output is obtained simply by taking the elementwise sigmoidal of $\mathbf{h}^{t}$. Notice that the $i$-th row of the matrix $\underline{\underline{M}}$ has zeroes in all columns corresponding to spins at distance larger than $\ell+1$ from the spin $i$, that is, the number of non-zero elements is constant with system size as long as $\ell$ is kept fixed. Therefore, generation can simply be implemented as a series of vector products, each considering only the nonzero elements of the matrix $\underline{\underline{M}}$. This allows one to compute the $i$-th element of vector $\mathbf{h}^{t}$ in a time $\mathcal{O}(1)$ in system size, therefore keeping generation procedures $\mathcal{O}(N)$. Moreover, the training time scales as $\mathcal{O}(N)$.
We point out that we can train directly the whole matrix $\underline{\underline{M}}$, with the additional constraint of putting to zero the elements corresponding to spins at distance larger than $\ell+1$. Conceptually, this approach can be interpreted as a reduced MADE, in which the weight matrix is not only triangular but has additional zeros. We will refer to this approach as $\ell$-MADE in the following. This method is easier to implement and can be faster in the training and generation processes, but has the disadvantage or requiring more parameters than the standard 4N approach, e.g. $\mathcal{O}(\ell^{d}N)$ for a $d-$dimensional lattice against the $(c+1)\ell N$ of 4N. Moreover, it loses the physical interpretation of 4N, as it can no longer be interpreted as a series of fields propagating on the lattice.
Wall-clock generation and training times are shown in Fig. S1. Comparison between Fig. 3b obtained using 4N and the analogue obtained using $\ell$-MADE is shown in Fig. S2.
Figure S1: Scatter plots of the generation (left) and training (right) elapsed real times at at different system sizes for $\ell$-MADE. Both training and generations have been performed using dummy data (and therefore $N$ needs not to be the square of an integer). For generation, we generate 10000 configurations of $N$ five times. For training, we train a model with $\ell=2$ for 40 epochs using an Adam optimizer once per each value of $N$. In both cases, elapsed times scale linearly with the system’s size, as further evidenced by the linear fits (black lines).
Figure S2: Ratio between the cross-entropy $|\langle\log P^{\ell}_{\text{NN}}\rangle_{\text{data}}|$ and the Gibbs-Boltzmann entropy $S_{\text{GB}}$ for different values of the temperature $T$ and of the parameter $\ell$ for 4N (a, left) and $\ell$-MADE (b, right). Both $|\langle\log P^{\ell}_{\text{NN}}\rangle_{\text{data}}|$ and $S_{\text{GB}}$ are averaged over 10 samples.
@@
Figure S3: Comparison of the probability distribution of the overlap obtained using different architectures: 4N with $\ell=10$ layers (12160 parameters), shallow MADE (32640 parameters), NADE with $N_{h}=64$ (32768 parameters) and TwoBo (4111 parameters). The black dashed line is obtained using data gathered via Parallel Tempering (PT), while solid color lines are obtained using the different architectures. From a quick qualitative analysis 4N performs comparably to MADE and NADE despite having less parameters. Moreover, it also performs better than TwoBo. Despite the latter having fewer parameters in this $N=256$ case, 4N has a better scaling when increasing the system size, $\mathcal{O}(N)$ compared to $\mathcal{O}(N^{\frac{3}{2}})$, as pointed out in the main text. The TwoBo implementation comes from the GitHub of the authors (https://github.com/cqsl/sparse-twobody-arnn) and was slightly modified to support supervised learning (which could be improved by a different choice of the hyperparameters used for training).
@@
Figure S4: Comparison of the probability distribution of the overlap obtained using different update sequences of the spins for $\ell=10$. The black dashed line is obtained using data gathered via Parallel Tempering (PT), while solid color lines are obtained using 4N with different sequence updates. Spiral: spins are update in a following a spiral going from the outside to the inside of the lattice; Raster: spins are updated row by row from left to right. It appears as the spiral update performs much better than the raster update for the 4N architecture in equal training conditions, something that is not observed when training, for instance, the MADE architecture. This effect could be related to the locality characteristics of the 4N architecture.
S4 Cross-entropy plots in lin-lin and lin-log scale
Figure S5: Absolute difference between $\langle\log P^{\ell}_{\text{NN}}\rangle_{\text{data}}$ at various $\ell$ and at $\ell=10$ for $T=1.02$ and different system’s sizes, in lin-lin and lin-log. Data are averaged over ten different instances of the disorder and the colored areas identify regions corresponding to plus or minus one standard deviation.